Find the indicated function values for the function g(x)=3x+3.




a. g(0) b. g(3) c. g(-2/3) d. g(-2b) e. g(b+5)

1
Expert's answer
2021-11-09T12:05:52-0500

a.

g(0)=3(0)+3

=0+3

=3


b.

g(3)=3(3)+3

=9+3

=12


c.

g(-2/3)=3(-2/3)+3

=-2+3

=1


d.

g(-2b)=3(-2b)+3

=-6b+3


e.

g(b+5)=3(b+5)+3

=3b+15+3

=3b+18